A body of mass 20kg falls from rest through a height of 50m and comes to rest having penetrated a distance of 6m into a sandyground.calculate the average force exacted into Sandy in bringing the body to rest taking g to be 9.8m/s^2
1
Expert's answer
2021-07-28T05:49:04-0400

The work-energy theorem says

"mgh=Fd"

Hence, the average force acting on a body in sandy ground

"F=\frac{mgh}{d}=\frac{20*9.8*50}{6}=1633\:\rm N"